OpenStudy (anonymous):

factor out this> x^4 -25 explain the answer if you could plz. thanks

OpenStudy (anonymous):

difference of squares a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

thanks . solve it plz

OpenStudy (anonymous):

Oh, I think you can do sqrt(x^4), and sqrt(25)

Parth (parthkohli):

@hal_stirrup Did you understand?

Parth (parthkohli):

Can you tell what \(\sqrt{25}\) and \(\sqrt{x^4}\) are?

OpenStudy (anonymous):